OpenStudy (anonymous):

The sum of two consecutive odd integers is 236. What is the smaller integer?

OpenStudy (anonymous):

236= x+1 + x+3

OpenStudy (anonymous):

x=116 so 116+1 117 :)

OpenStudy (anonymous):

ok i was gonna work it out but thanx

OpenStudy (anonymous):

you can still work it out :P

OpenStudy (anonymous):

thanx

OpenStudy (apoorvk):

The form "2n+1" is always an odd no, since it's a '1' added to an even no (i.e. 2n). So, 'et the smaller of the two odd integers be '2n+1'. So the next integer would be '2n+3'. Now, their sum = (2n+1) + (2n+3) = 236 Solve this to find out 'x'! Enjoy!

OpenStudy (apoorvk):

*find out 'n'. And then calculate "2n+1" for the smallest one. @timo86m I think it's always safe to assume a "2n+1" form for an odd no. what do you say?

OpenStudy (anonymous):

well 117+119 works 2 consecutive odd integers that add up to 236 it works.

OpenStudy (anonymous):

(2n+1) + (2n+3) = 236 i solve for n i got 58? then what?

OpenStudy (anonymous):

2 + (x + y) = (2 + x) + y is an example of which algebraic property?

OpenStudy (apoorvk):

After that, we find out the smallest knowing that it is is "2n+1", so it's 2(58)+1 = 117. Same as what you got. Similarly for the larger no. Well your method does work in this case, but may just turn out a bit troublesome if there's something involved like a negative and positive integer. 'May' just..

OpenStudy (apoorvk):

@terrell352 that's called a associative property. An associative property for any operation ' * ' is applicable whenver a * (b * c) = (a * b) * c Also, please post each new question in a new thread. Thanks!

OpenStudy (apoorvk):

*whenever